Instrument Log

60 MT T/H "A" (Southeast) Side C1 Lamont, OK

Date

Technician

Comments

6/20/16

JEM

Maintenance was being performed on the 25M T/H from 1715 GMT to 1730 GMT. Data may be effected during this time. RTS @1730 GMT.

6/14/16 6/15/16

CAW/TMG

SE tower booms needed to be lowered to install new cameras on tower. Lowered SE booms at 1800 GMT on 6/14. Booms remained down overnight while working on camera mounts. Raised booms after cameras were installed and RTS at 2025 GMT on 6/15.

12/3/15

JEM

Instrument will be effected while maintenance is being performed on the 25M MFR. Effected time period is 1520 GMT to 1605 GMT.

10/27/15

JEM

The US and UIR needs to be changed for calibration. The SE boom was lowered @1522 GMT for swapping the UIR and US. 605M T/H data will be effected during this time. Raised the boom and RTS @1620 GMT.

10/26/15

CAM

D. Cook

Instrument mentor on site for annual 25 and 60M T/H calibration change outs. Lowered the SE booms at 1435 GMT for maintenance. R/R 60M *T/H sensor. Raised booms and RTS at 1949 GMT.

*Old S4520023 *New S4930004

8/25/15

JEM

DQPR #4842 indicates that the battery voltage has started to drop out periodically. Battery voltage observed at 12.2vdc. Troubleshooting found the battery and charger were pretty warm and the battery terminals were also heavily corroded. Replaced the battery and power supply @1405 GMT. Battery voltage increased to approximately 12.7vdc. RTS @1410 GMT.

2/2/15

JEM

The 25M MFR needs to be reinstalled. 60M T/H data will be affected from 2048 GMT to 2100 GMT. RTS @2100 GMT.

1/30/15

JEM

The boom was lowered for 25M MFR maintenance. Data will be affected from 1932 GMT to 1952 GMT. RTS @1952 GMT.

1/27/15

JEM

The boom was lowered for MFR25M maintenance. Data will be affected from 2105 GMT to 2120 GMT.

10/28/14

CAM

D. Cook

Instrument mentor on site for annual 25M and 60M T/H calibration change outs. Lowered the SE booms at 1520 GMT for maintenance. R/R 60M *T/H sensor. 60M ventilator fan had failed and was not operational. R/R ventilator fan. Raised booms and RTS at 2230 GMT.

*Old S4930004 *New S4520023

8/29/14

JEM

The TWR10X has not collected since the 60M tower was lowered on 8/27/14. Troubleshooting found that the ribbon cable had worked itself loose from the SC32A. Reconnected cable and installed keeper screw. Collections are now successful. No data will be lost. RTS @1745 GMT.

8/27/14

JEM

The 25M US needs to be replaced for annual calibration. The East boom was lowered from 1330 GMT to 1348 GMT to replace the 25M US. Data will be effected during this period. RTS @1350 GMT.

6/23/14

CAM

No data had been collected from 60 T/H since 1208 GMT on 6/21. Found the UPS that powers the comms equipment in the AOS trailer was down. Power cycled UPS and restored power at 1220 GMT. Normal data collections resumed at next interval.

5/27/14

CAM

No data had been collected from instrument since 1205 GMT on 5/25. Found the UPS that powers the comms equipment in the AOS trailer was down. Breaker for UPS supply line was tripped. Reset breaker at 1325 GMT. UPS powered up and comms equipment power restored. Normal data collections resumed at next interval.

4/30/13

JEM

Received an email from the mentor with the following request: Please replace the blue resistor with the brown resistor for the tower SE side 60 m (upper part of panel, differential channel 4). Swapped the resistors @2020 GMT. RTs @2030 GMT.

4/24/13

Twr Sys

Tower Systems on site for tower maintenance. Unsnagged SW winch cable from foot pegs and straightened bent sections of boom guide rails. Adjusted boom vertically on winch cable for proper alignment with top connectors. Winch cable will need to be remarked next time the booms are completely lowered.

2/27/13

CAM

Cycled AC power to southeast 60M level at 1921 GMT to reset network camera. No problem with T/H instrument.

2/26/13

CAM

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm. Power was restored at 1920 GMT. Instrument RTS status O.K and data collections resumed once SDS operations were restored.

2/8/13

JEM

DQPR #3616 requests the 25M IRT lens be replaced. Lowered the SE booms @1942 GMT. R/R the IRT lens @1948 GMT. Raised the SE booms and RTS @1957 GMT.

12/24/12

JEM

Tower booms need to be lowered to clean the 25M IRT len and verify IRT settings. Tower was lowered @1544 GMT. Raised the tower @1627 GMT. Status OK. RTS @1630 GMT.

11/20/12

CAM

Local electrical service was shut off at 1410 GMT so that Kay Electric could upgrade its power lines from 7.2 KV to 14.4 KV in Grant County. Power was restored at 2115 GMT after electrical upgrades were complete.

11/14/12

JEM

Mentor requests a change in the data logger program. Changed statement 1, parameter 8 from 1.0011 to 1.0000. Also verified statement 3, parameter 8 as being 1.0000. Recompiled the program @1614 GMT. Verified readings and RTS @ 1615 GMT. Notified mentor of the change.

11/13/12

CAM

Site power was partially shut down at 1400 GMT in order to replace the 7 transformers that service the buildings and all instruments south of the main compound. Power was restored at approximately 1930 GMT after all transformers were replaced. Tower instruments resumed normal operation.

11/5/12

CAW

Lowered booms at 1338 GMT for 25 MFR 90-day calibration. Raised booms and RTS at 1349 GMT after MFR head removed. Lowered booms again for reinstallation at 2025 GMT. Raised booms and RTS at 2036 GMT.

10/31/12

JEM

The boom needs to be lowered to verify the 25M IRT serial number. Booms were lowered form 1513 GMT to 1535 GMT.

10/9/12

CAM

D. Cook

Instrument mentor on site for annual 25 and 60M T/H calibration changeouts. Lowered the SE booms at 1900 GMT for maintenance. R/R 60M *T/H sensor. Raised booms and RTS at 2145 GMT.

*Old S4520023 *New S4930004

9/20/12

PAD

CF experienced a power outage on 9/19/12 at 1528 GMT. Last SDS collection for twr10x for side A of tower was on 9/19/12 at 1500 GMT. SDS computer is able to connect with CDL but communications are lost shortly after connection. Connected to CDL directly with laptop. Values are shown but datalogger freezes every 10-15 seconds. Any commands given to CDL will result in loss of connection. Unable to retrieve program from storage module. R/R CDL but was still unable to retrieve program. Took storage module to electronics lab. Technician was able to connect with CDL long enough to copy program into another place on storage module. Program was then retrieved and loaded onto replacement CDL. RTS at 1940 GMT. CDL Model 10X OLD SN 2966. New 25892.

7/13/2012

JEM

The SE tower elevator was lowered to perform work on the 25M US. Lowered the SE tower elevator @1943 GMT and raised it @ 2002 GMT.

6/26/12

CAM/PAD

SDS is not collecting data from 25M IRT. Lowered side A booms at 1837 GMT. Raised booms at 2007 GMT. No issues know with instrument at this time. RTS at 2007 GMT.

6/4/12

JEM

Problems continue with the 60M T/H reading low. The problem has gotten worse and it is reading nearly 40C lower than other sensors. Lowered the tower @1410 GMT. Extensive troubleshooting and testing revealed multiple problems that were likely contributing to the problem. All sensor connections on the boom were found tight but were loosened and re-seated anyway. This brought the temperature back up to approximately 2~3C lower than the 25M sensor. Further checks found a bad leg on a 100 Ohm resistor that was across 1H and 1L on the wiring panel. Replaced this resistor with a new one due to the length of the remaining leg on the old resistor. Also found that the ventilator on the 60M T/H was not receiving proper voltage and therefore was only drawing air at 1.7 m/s while the 25M T/H ventilator was drawing air at 3.1 m/s. R/R 60M T/H ventilator power supply. Both ventilators are now drawing air at 3.1 m/s. While at ground level the 60M and 25M T/H were within 0.2C and 1% of each other. Raised the tower @1637 GMT. After raising the tower all temperature readings were within 1C of each other. All humidity readings were within 4% of each other. RTS @1640 GMT.

6/1/12

JEM

DQPR #3407 request the 60M T/RH sensor be replaced. Lowered the SE booms at 1412 GMT. Replaced *T/RH sensor. Raised the SE booms and RTS @1512 GMT. *Old-S4520022 *New-S4520023

5/24/12

PAD

DQPR 3407 states there is a low bias for 60 T/H probe on side A of tower. Mentor suspects bad connection and suggests lowering/raising tower. Lowered side A 25 and 60m booms at 2115 GMT. All indicator lights for signal and AC were lit after raising booms. RTS at 2130 GMT.

5/8/12

CAW

Lowered booms at 1258 GMT for 25 MFR 90-day calibration. Raised booms and RTS at 1310 GMT after MFR head removed. Lowered booms again for reinstallation at 1522 GMT. Raised booms and RTS at 1535 GMT.

3/8/12

PAD

Radon Gas Analyzer inlet pipe has separated approximately 40M above the ground. Contractor is on site today to climb tower and repair inlet tube. Power to tower needs to be shut down before contractor can climb. Shut down power at base of tower to 25 and 60M levels, side B at 2023 GMT. After repairs were made and complete inspection of inlet pipe was performed, power was restored. No issues with this instrument are known at this time. RTS at 2257 GMT.

2/8/12

JEM

The 25M MFR is scheduled for an upgrade. Tower was lowered from 1855 GMT to 2212 GMT. 60M T/H will be affected during this time.

11/1/11

CAM

D. Cook

Instrument mentor on site for annual 25 and 60M T/H calibration changeouts. Lowered the SE booms at 1815 GMT for maintenance. Replaced T/H sensor but new sensor (SN S4930005) was defective with temperature reading of -9999.9. Placed original sensor back into service. Raised booms and RTS at 2130 GMT. *SN S4520022

7/19/11

CAW

25M MFR is due for calibration. Side A tower booms were lowered at 1910 GMT to remove 25M MFR. Tower was raised at 2107 GMT. No issues known with this instrument at this time. RTS at 2110 GMT.

4/15/11

JEM

The instrument has not collected since the 0400 hour. Due to other instruments that are down the source of the problems seems to be the Rocketport hub. The Rocketport power light is flashing. Power cycling the Rocketport hub did not correct the issue. Notified the IT department of the issue.

4/7/11

CAW

MFR is due for calibration. Lowered tower at 1200 GMT and raised tower at 1215 GMT. Calibrated instrument. Lowered tower at 1433 GMT, installed instrument, and raised tower at 1444 GMT. Instrument only affected by lowering tower to calibrate 25M MFR.

11/23/10

CAW

Lowered the SE booms at 1825 GMT for 25MFR calibration. Raised booms and RTS at 2010 GMT. No problem with this instrument.

10/19/10

CAM

Tower Systems on site for tower inspections. Need to secure power for maintenance. Secured AC power to 25m and 60m levels at 1407 GMT. Restored power following inspection at 1620 GMT.

10/18/10

CAM

D. Cook

Instrument mentor on site for annual 25 and 60M T/H calibration changeouts. Lowered the SE booms at 1445 GMT for maintenance. R/R 60M *T/H sensor. Raised booms and RTS at 2013 GMT.

*Old S2920010 *New S4520022

8/19/2010

JEM

The 60T/H was disrupted during maintenance on the 25MFR from 1430 GMT to 1500 GMT.

8/18/2010

JEM

The 60T/H was disrupted during maintenance on the 25MFR from 1826 GMT to 1841 GMT.

7/14/10

CAM/JEM

60M tower SE booms need to be lowered to remove GPS array repeater. Lowered SE booms at 1445 GMT. Removed GPS equipment. Raised booms and RTS at 1556 GMT.

6/11/10

PAD

60M tower cameras are not working. Computer department requests both sides of tower power cycled at the 60M level to reset cameras. Power cycled Side A and Side B at 60M level at 1738. No problem with this instrument known at this time. RTS at 1753 GMT.

5/11/10

CAW

No problem with instrument at this time. 25M MFR is due for calibration. Side A tower was lowered at 1848 GMT to remove 25M MFR and raised at 1900 GMT. Tower was lowered again at 2040 GMT to install 25M MFR and raised at 2052 GMT. RTS at 2055 GMT.

5/3/10 

JEM

The mentor requests the CR10X be replaced.  Powered the logger off at 1436 GMT. R/R the *CR10X data logger @1437 GMT. Powered the logger back on and loaded the program at 1438 GMT. Reading looked good after the program started running. RTS @1450 GMT.

*Old 4503 *New - 44204

2/23/10

PAD

Last SDS collection was at 1009 GMT. Power cycled 60 T/H at 1608 GMT. SDS successfully collected data at 1709 GMT. RTS at 1623 GMT.

2/16/10

CAM

25M MFR needs to be removed for calibration. There is no problem with the SE T/H sensors. Lowered SE booms at 1455 GMT and removed MFR. Raised booms at 1507 GMT. Lowered booms again at 1712 GMT to reinstall MFR after it was calibrated. Raised booms and RTS at 1722 GMT.

11/19/09

CAM

SDS instrument status page indicated no data had been collected since 2105 GMT on 11/18. T/S found the UPS in the AOS trailer which powers the tower communications equipment had no output. AC LINE and ALARM LEDs were lit on the UPS. Power cycled UPS at 1342 GMT and power was restored. RTS at 1345 GMT.

11/9/09

CAM

25M MFR needs to be removed for calibration. Lowered SE booms at 1405 GMT and removed MFR. Raised booms at 1414 GMT. Lowered booms again at 1623 GMT to reinstall MFR after it was calibrated. Raised booms and RTS at 1633 GMT.

9/21/09

D. Cook

CAM

Instrument mentor on site for annual 25 and 60M T/H calibration changeouts. Lowered the SE booms at 1450 GMT for maintenance. R/R 60M *T/H sensor. Raised booms and RTS at 1735 GMT.

*Old S2920011 (HMP35D) *New S2920010 (HMP35D)

8/25/09

CAW

SE booms lowered from 1320 to 1330 GMT to remove MFR sensor for maintenance.

5/19/09

CAM

SE booms lowered from 1743 GMT to 1758 GMT for 25m MFR maintenance. No problem with this instrument.

5/13/09

CAM

25M MFR needs to be reinstalled after it was calibrated. Lowered booms at 2146 GMT and installed MFR. Raised booms and RTS at 2201 GMT.

5/13/09

CAM

25M MFR needs to be removed for calibration. Lowered booms at 1431 GMT and removed MFR. Raised booms and RTS at 1446 GMT.

10/1/08

CAM

Lowered SE booms at 1850 GMT to clean IRT lens. Raised booms and RTS at 1905 GMT.

6/18/08

CM

CF experienced power outages on 6/16/08 at 0429 GMT and again at 1834 GMT. SDS last collected SE boom 60M T/H data on 6/16/08 at 1008 GMT. Troubleshooting found SDS is communicating with datalogger. Suspect a hung or corrupt file. At 1445 GMT, connected datalogger with laptop and collected all data. Power cycled datalogger at 1510 GMT. This did not fix collection problem. At 1830 GMT, collected all data and retrieved DLD program for backup. Reset intermediate and final storage by entering same size memory number. This cleared datalogger memory without altering program. SDS successfully collected data at 1908 GMT. RTS at 1910 GMT.

3/10/08

PAD

GPS array mentor request repeater located on side A of tower power cycled. At 1427 GMT, power cycled side A 60M level. No problems known with this instrument at this time. RTS at 1442 GMT.

7/26/07

D. Cook

CAM

Instrument mentor on site for annual 25 and 60M T/H calibration changeouts. Lowered the SE booms at 1333 GMT for maintenance. R/R 60M *T/H sensor. Raised booms and RTS at 1920 GMT.

*Old S4520015 (HMP45D) *New S2920011 (HMP35D)

10/11/06

D. Cook

CAM

Instrument mentor on site for annual 25 and 60M T/H calibration changeouts. Lowered the SE booms at 1447 GMT for maintenance. R/R 60M *T/H sensor. Raised booms and RTS at 2141 GMT. *Old S2920011 (HMP35D) *New S4520015 (HMP45D)

5/31/06

PAD

Jane McKinney request 60m tower lowered to record model and serial numbers from the wireless data transceiver located on SE 60M boom. Lowered side A of tower at 1755 GMT. Raised tower at 1851 GMT. No problem with this instrument. RTS at 1855 GMT.

11/30/05

PAD

Mentor request data logger battery changed out with a new one. At 2145 GMT, R/R data logger battery. RTS at 2200 GMT.

11/28/05

PAD

Last SDS collection was at 0502 GMT. Found data logger power supply wire had been cut by rodents. Removed power supply at 1550 GMT. Mark Klassen repaired wire. Installed power supply and protective sheath. RTS at 1650 GMT.

8/5/05

D. Cook

CAM

Instrument mentor back on site to finish the annual 25 and 60M T/H calibration checks. Lowered the SE booms at 1343 GMT for maintenance. Completed calibration checks. Raised booms and RTS at 1632 GMT.

8/4/05

D. Cook

GRS

SE booms need to be lowered for 25m and 60m T/H calibrations. Tower Systems techs also on site for tower inspection. Turned the 25 and 60M levels power off at 1330 GMT so Tower Systems could service tower. Lowered the SE booms at 1828 GMT for maintenance. R/R 60M *T/H sensor. Raised booms and RTS at 2016 GMT. *Old - S4520015 (HMP45D) *New - S2920011 (HMP35D)

6/10/05

MR

The SE tower needs to be lowered due to maintenance on other equipment. Lowered the tower at 1945 GMT. Raised the tower at 2045 GMT. RTS at 2100 GMT.

5/25/05

MR

The maintenance on the PGS equipment is finished and the SE elevators need to be raised. Raised the SE elevators. RTS at 2145 GMT.

5/24/05

MR

The SE elevators need to be lowered due to maintenance on the PGS equipment. Lowered the SE elevators at 1410 GMT. These will be left down overnight due to maintenance issues on the PGS equipment.

9/27/04

D. Cook

CAM

Instrument mentor on site for annual 25 and 60M T/H calibration changeouts. Lowered the SE booms at 1438 GMT for maintenance. R/R 60M *T/H sensor. Raised booms and RTS at 2056 GMT. *Old S2920009 (HMP35D) *New

8/13/04

PAD

During troubleshooting GPS array system problems, 60M SE boom ,(Side A), needed to be power cycled . Power cycled 60M instruments on SE boom at 1603 GMT. RTS at 1616 GMT.

8/5/04

MR

The SE elevator needs to be lowered due to work on other instruments. Lowered the tower at 1410 GMT. Raised tower at 1445 GMT. RTS at 1450 GMT.

7/21/04

MR

The tower South east needs to be lowered and the instruments need to be cleaned due to the farmer working the ground under the tower. Lowered the elevator at 1526 GMT. Cleaned the sensor cap and ventilator at 1610 GMT. Raised the elevator at 1615 GMT. RTS at 1626 GMT.

7/1/04

PAD

60M T/H last successful collection was on 6/30/04 at 1602 GMT. Power cycled data logger at 1626 GMT. RTS at 1641 GMT.

4/15/04

CM/PD

SE booms need to be lowered for 25M MFR maintenance. Lowered booms at 2104 GMT. Raised booms and RTS at 2151 GMT.

2/23/04

MR

The SE tower connections need to be reseated for other equipment. Lowered the elevator about 1 foot and the raised the elevator. RTS at 1853 GMT.

2/20/04

MR

The SE elevator needs to be lowered due to work on other equipment. Lowered the elevator at 1445 GMT. Raised the elevator at 1515 GMT. RTS at 1515 GMT.

2/13/04

MR

The elevator needs to be lowered due to maintenance on other equipment. Lowered the elevator at 1752 GMT. Raised the elevator at 2134 GMT. RTS at 2140 GMT.

2/10/04

MR

T/S found that the T/H sensor needs to be replaced. Lowered the booms at 1905 GMT. R/R *T/H sensor. Raised the booms at 2315 GMT. RTS at 2330 GMT. *Old - S4930004 *New - S2920009

1/29/04

PAD

Instruments on 60M West boom were power cycled due to problems with 60M CO2 flux. Power cycled 60M instruments at 2209 GMT. RTS at 2302 GMT. No problems with this instrument.

12/8/03

MR

The tower needs to be lowered due to maintenance on other equipment. Lowered the tower at 1940. Raised the tower at 2105. RTS at 2105.

11/19/03

MR

The 25M and 60 M tower elevator needs to be lowered due to maintenance on other equipment. Lowered the SE tower at 2105. Raised the tower at 2120 GMT.

11/10/03

MR

The instrument mentor wants the data logger changed out. R/R *data logger. Downloaded the program into the data logger. RTS at 2105. *Old - X2966 *New - X4503

10/22/03

MR

The instrument mentor wants the lightening protection bypassed in order to help isolate some data problems. Removed the lightening protection. RTS at 1845 GMT.

10/7/03

MR

The tower connections need to be reset. Lowered the tower about 3 inches at 1530. Raised the tower at 1531. RTS at 1545.

10/6/03

MR

DQPR #118 requests that the 25M and 60M towers be lowered and data logger voltages be checked. Lowered the towers at 1713 GMT. Connected the data and power jumper cables up at 1724. Allowed the system to stabilize from 1724 until 1800 GMT. Measured the required voltages as follows; AG-G=0.0, 3H-AG=.769, 3H-G=.769, 3L-AG=0.007, 3L-G=0.007, 6H-AG=.727, 6H-G=.727, 6L-AG=0.004, 6L-G=0.004. Raised the towers back up at 1826. RTS at 1828. Notified the mentor of the readings.

10/6/03

MR

The instrument mentor wants the internal battery checked on the data logger for the 60M SE sensors. Checked the battery. The reading is 3.12 volts. RTS at 1740 GMT.

9/24/03

MR

The instrument mentor wants some of the voltages measured and recorded. Measured the following votlages at the following locations. G to AG = 0.0 v, 3H to AG = .659 v, 3H to G = .659, 3L to AG = .008 v , 3L to G = .008 v, 6H to AG = .613, 6H to G = .613, 6L to AG = .004, 6L to G = .004. RTS at 1510 GMT.

9/11/03

MR

DQPR # 118 requested that the wiring for the 60 M T/H south side be checked. Checked the wiring on the data logger as requested in the DQPR. The 3L and 6L data logger channels did not have a jumper to AG. Installed the jumper to AG from 3L and 6L as requested. Notified the mentor. RTS at 1405.

8/19/03

MR

The elevator needs to be lowered due to maintenance on other equipment. Lowered the tower at 1415 GMT. Raised the tower at 1603. RTS at 1615.

8/15/03

CAM

Dave Cook on site for annual 25 and 60M T/H calibration changeouts. Lowered the SE booms at 1341 GMT for maintenance. R/R 60M *T/H sensor. Raised booms and RTS at 1920 GMT. *Old - S1720003 *New S4930004

8/14/03

CM

SE booms lowered a few inches at 1900 GMT to tension the cable between the ground and 25M boom. Raised booms at 1914 GMT. Only the 60M boom made connection at that time due to the added tension. Lowered booms again at 1925 GMT and repositioned the 25M boom higher on the cable. Raised booms and RTS both 25 and 60M levels at 1935 GMT.

8/14/03

CM

SE booms lowered for tower system maintenance. There was no failure with this instrument. Lowered boom at 1443 GMT. Raised boom and RTS at 1515 GMT.

8/5/03

MR

The elevator tower needs to be lowered due to maintenance on other equipment. Lowered the tower at 1836 GMT. Raised the tower at 1910 GMT. RTS at 1921 GMT.

3/21/03

MR

The elevator needs to be lowered due to maintenance on other equipment. Lowered the tower at 1824 GMT. Raised the tower at 2024 GMT. RTS at 2024 GMT.

3/18/03

JEM

The 60M T/H has not collected since 3/17/03 @1502 GMT. The power strip that powers the bit driver in the AOS trailer was found powered off. Powered the power strip back on @1505 GMT. The power strip was likely powered off by accident while rearranging the AOS system. RTS @1515 GMT.

3/17/03

MR

The tower will be lowered due to maintenance on other equipment. Lowered the tower between 1625 and 1705. RTS at 1705.

2/21/03

MR

Due to maintenance on other equipment the elevator needs to be lowered. Lowered the elevator at 1953 GMT. Raised the elevator at 2056 GMT. RTS at 2056 GMT.

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT. Power was restored to the Central Facility at approximately 1945 GMT. RTS @2000 GMT.

1/30/03

JEM

The 60M Southeast temperature has been too high. Lowered Southeast 25M and 60M booms @1655 GMT. The 60M T/H Aspirator fan was not working. R/R Aspirator fan @1715 GMT. The new fan would not work either. Upon checking the AC/DC power supply, it was found to be unplugged. Plugged the AC/DC power supply in and the fan began working. Removed the new aspirator and reinstalled the old aspirator fan @1740 GMT. It is possible, but unlikely that the power supply vibrated out of the receptacle. The power supply was likely left unplugged. Placed a stiff piece of foam behind the power supply to ensure that it could not vibrate out. At 1750 GMT the 60M temperature was reading 5.5 degrees and the 25M temperature was reading 5.7 degrees. Raised the 60M and 25M booms back into place @1759 GMT. RTS @1810 GMT.

12/16/02

MR

The tower needs to be lowered due to maintenance on other equipment. Lowered the tower at 1639 GMT. Raised the tower at 1705 GMT. RTS at 1710 GMT.

11/7/02

CM/MR

SE booms lowered from 1557 GMT to 1755 GMT to troubleshoot 25M IRT communications problems.

11/6/02

CM/MR

SE booms lowered from 2017 GMT to 2222 GMT to troubleshoot 25M IRT communications problems.

10/31/02

CAM

Lowered the SE booms at 1445 GMT for continued T/S of 9999s in the 60M T/H data and to replace sensor if necessary. Verified that all color coded wiring was connected properly, which it was, but did find a jumper wire laying in the bottom of the 60M junction box that was not being used. Opened the 25M junction box for comparison and observed a jumper on the terminal strip between the violet (gnd) and red (sensor gnd) wires. Inserted the 60M jumper in the same location as the 25M and the 60M sensor then reported good data. Note there is no red wire on the new sensor that was installed at the 60M level on 10/25, whereas the old sensor did have a red wire. Raised the booms and RTS at 1618 GMT. Monitored locations 1-6 from the ground and confirmed the readings were still good.

10/29/02

CAM

Quicklook data plots indicate that the SE 25 and 60M T/H sensors are reading 9999s or 0s. Monitored CR10X locations 1-6 with the handheld keypad and confirmed all were reporting bad values. Checked location 10 (battery voltage) and it too was reporting a bogus value of 3, even though the battery voltage measured 13V with the DVM. Suspected the program may be corrupted. Reloaded the program with the same results. Tried a new datalogger with the same results (returned original datalogger to service). Replaced the CR10X *wiring panel but still no change in status. Lowered the SE booms at 2135 GMT to check at ground level. Connected only the 25M jumper cables and all 25M values reported good. Connected the 60M DC signal jumper cable and all values at both the 25M and 60M immediately went bad. Verified that the 60M T/H sensor connections were correct. Disconnected the 60M T/H sensor in the junction box and the 25M values returned good. Suspect the 60M T/H sensor is bad. Left the 60M T/H sensor disconnected and raised the SE booms at 2218 GMT. The 25M readings are now reporting good data with the 60M sensor disconnected. Will contact instrument mentor on 10/30 to see which sensor to install at the 60M level. RTS at 2220 GMT with only the 25M T/H operational. *Old 1375 *New - 17803

10/28/02

CAM

The SE 25 and 60M T/H sensors are reading 9999s or 0s. Lowered the SE booms slightly at 1731 GMT and then raised to reseat connections. RTS at 1733 GMT. Readings are still bad. Due to approaching rain, will lower booms fully to T/S tomorrow.

10/25/02

MR

The instrument mentor want the 60 M T/H probe replaced. R/R the probe. RTS at 1920 GMT. **Old Sn# S4930004 New Sn# S1720003**.

10/24/02

MR

The tower needs to be lowered due to maintenance on other equipment. Lowered the tower at 1415 GMT. Raised the tower at 1430 GMT.

9/23/02

MR

The elevator needs to be lowered due to work being performed on the 25M MFRSR. Lowered the tower at 1930 GMT. Raised the tower at 1943 GMT. RTS at 1945 GMT.

9/10/02

MR

The instrument mentor wants the tower lowered so work can be done on the PGS sampling system. Lowered the tower at 1733 GMT. Raised tower at 1837 GMT. Status ok. RTS at 1848.

9/3/02

MR

The 60M elevator needs to be lowered to check an instrument for the PGS system. Lowered the tower at 1940 GMT. Raised the tower at 2030. RTS at 2040.

8/2/02

D. Cook

The 25M T/H was not operating in a normal manner. T/S found that the CR10X was the problem. R/R the *CR10X @ 1400 GMT. The program uploaded from the storage module. Set the data logger clock. Status OK. RTS 1430 GMT. *Old - X4503 *New - X2966.

8/2/02

D. Cook

Booms lowered from 1354 to 1912 GMT for tower maintenance and 25/60m T/H calibrations. R/R *T/H sensor for annual calibration. *Old - S1720003 *New S4930004

8/1/02

D. Cook

TSI on site. Booms lowered from 1300 to 2130 GMT for tower maintenance.

4/11/02

CAM/MR

Lowered the SE booms at 2003 GMT for 25M MFR maintenance. Raised the booms and RTS at 2106 GMT.

4/2/02

CAM

The SE booms needed to be lowered momentarily to reseat the MFR connections. Lowered the SE booms at 1754 GMT. Raised the booms and RTS at 1755 GMT.

2/13/02

MR

The tower needs to be lowered for maintenance on other equipment. Lowered the tower at 1500. Raised the tower at 1505. RTS at 1515 GMT.

12/19/01

MR

The tower needs to be lowered due to calibration on the PGS system. Lowered the tower at 2055 GMT. Raised the tower at 2110 GMT.

11/01/01

MR

Vic Morris requested that the upwelling wiring be checked due to the readings on the US and UIR are bad. Lowered the tower at 2050 GMT. T/S found that the 100K precision resistor between the CR10 and multiplexer was not installed. Installed the resistor and notified the mentor of the change. Raised the tower at 2103 GMT.

10/4/01

MR

Joe Michalsky wants to run the calibrations on the 25M MFR a second time. Lowered the east side of the tower at 1828 GMT. Raised the tower at 1900. RTS at 1905 GMT.

10/3/01

MR

Lowered the east side of the tower at 2006 GMT for 25M MFR calibration. Raised the east side of the tower at 2057. RTS at 2105.

10/2/01

MR

The tower had to be lowered due to other maintenance. Lowered the tower at 1548. Raised the tower at 1600. RTS at 1603.

9/27/01

MR

Lowered the 25 and 60 meter tower due to maintenance being performed on the 25 meter tower. Lowered the tower at 1503. Raised the tower at 1912.

8/15/01

CAW

SE booms lowered from 1530 to 1555 GMT to R/R 25M US radiometer.

8/7/01-8/8/01

MR

The instrument mentor is on site to replace and calibrate the 25M and 60M T/H probes. Lowered the southeast towers at 1918. The mentor R/R the 60M T/H probes. The mentor discovered that the new 60M probe was reading bad but due to time constraints we were unable to replace the probes and raise the tower. Left the tower down for the evening and the mentor will replace the probe on the morning of 8/8/01. On 8/8/01, at 1330 GMT, the mentor R/R the probe and performed calibrations. The tower was then raised on 8/8/01 at 1840 GMT. *Old - S4520023 (sensor that was installed, then replaced the next day - S1720005) *New - S1720003

8/7/01

MR

Tower systems is on site to perform maintenance on the 25M elevator and the 60M tower. Lowered the southeast 25M elevator. Tower systems replaced one of the connectors on the tower and performed general maintenance on the elevator and tower components. RTS at 1910 GMT.

7/13/01

MR

The 25m tower was lowered due to maintenance on the 60M PGS system. Lowered the tower at 2000 GMT. Marc Fisher performed maintenance on the PGS system. Raised the tower at 2130. RTS at 2145 GMT.

7/13/01

MR

The SDS is unable to talk to the 25M T/H and 60M T/H southeast. Lowered the tower to power cycle the instruments at 1440 GMT. Raised the tower at 1441 GMT. Communications were restored.

7/12/01

MR

The 60m tower was lowered due to maintenance on the 60M PGS system. Lowered the tower at 2100 GMT. Marc Fisher performed maintenance on the PGS system. Raised the tower at 2135. RTS at 2145 GMT.

11/27/00

CAW

Lowered the SE 25m and 60m booms from 1840 to 2035 GMT to clean the chilled mirrors.

9/22/2000

MR

The 25 meter and the 60 meter humidity readings are reading -9999. Checked the connections on the wiring panel and they checked O.K. Checked the connection inside the tower connection box and they checked O.K. Reloaded the data logger program but this did not fix the problem. Changed out the *data logger and the readings returned to normal. Downloaded all the data off the storage module and cleared it. *Old - X08317 *New - X4503

9/21/2000

CAM

The site scientist advised that the 25 and 60 meter humidity levels on the SE side (CR 10X) were reading approximately 50% high since 9/19/00. Instrument mentor recommended power cycling the data logger first and then lowering and reseating the booms if the power cycle did not correct the problem. Checked the humidity readings at 1845 GMT on 9/21 and found they were both reading in the low 80% range. The actual humidity was approximately 30%. Power cycled the CR 10X data logger at 1845 GMT. This did not correct the readings. In fact, the readings then went to -99999 for locations 3,4,5,6, 13 and 14 on the datalogger. Lowered the booms slightly and then reseated at 1850 GMT. This also produced no change in the readings (still -99999). Power cycled the data logger again at 1852 GMT with no change. Checked the datalogger program and found it to be OK. Disconnected the sensors and measured the output directly to be 314mV, which is accurate. Connected a spare humidity sensor to the datalogger at ground level and found that it too read -99999. The datalogger needs to be replaced. RTS at 2030 GMT and datalogger will be replaced tomorrow.

8/14/2000

CAM/DC

David Cook on site for annual sensor calibrations and change outs. Lowered all tower booms at 1447 GMT. Connected power and data jumper cables and let sensors stabilize at ground level for 3 hours and checked sensor performance. R/R the east 60M *temperature/humidity sensor for annual calibration. Raised the booms and RTS at 2138 GMT. *Old - S1720005 (model HMP35D) *New - S4520023 (model HMP45D)

1/6/2000

CM/CW

25M and 60M booms lowered from 1941 GMT to 2123 GMT for installation of new GPS antenna.

11/16/99

CM/SS

The A side booms were lowered from 1700 GMT to 1820 GMT to finish insulating the 25M IRT.

11/15/99

CM/GS

The A side booms were lowered from 2011 GMT to 2217 GMT to insulate the 25M IRT.

9/3/99

RS

The 60 meter T/H sensor will be with out power for a period of two hour starting at 1800 till 2000 and then on for for two hour at 2000 till 2200 then off for two hour hours 2200 till 2400 and then back on. This is being performed by request of the 915 RWP mentor to see if the GPS repeater is causing interference with the 915 RWP. Turned the power on/off to the 60 meter position on the "A" side of the tower as requested by the 915 RWP mentor.

5/19/99

5/20/99

CAM/DC

David Cook on site for annual sensor calibrations. Lowered the 60M A side boom at 1755 GMT on 5/19. R/R the *temperature/humidity sensor for calibration. Performed ice bath check on new temperature sensor and made calibration factor adjustment as necessary. Sensor then checked out O.K. Left the boom at ground level overnight for ECOR maintenance. Raised A side boom and RTS status O.K. at 1915 GMT on 5/20/99. *Old - S1720002 (Model HMP35D) *New - S1720005 (Model HMP35D).

12/8/98

CM/GS

The 25m boom (side A) was lowered at 2105 GMT for IRT maintenance. This resulted in the 60m T/H being taken off line as well. Raised the booms and RTS at 2153 GMT.

9/18/98

CM/MW

SDS personnel advised that no data had been collected from 25m or 60m 10X T/H since 8/26/98. The collection status showed successful since that time but no files were being ingested. Checked CR10X data logger status using CR10KD handheld and found it to be operational. Attempted to collect data from the storage module but could not establish a link. Attempted to monitor data logger using laptop computer but only received garbled characters back. Disconnected storage module from data logger and could then successfully monitor logger. Removed storage module to the lab and connected to test system. Was able to link with storage module in lab but module contained no data. Erased and reset storage module and placed back in system. Still could not link with storage module at tower nor could the data logger be monitored while the storage module was connected. Removed both the data logger and storage module to the lab at 1745 GMT and placed in test system. Using the same cables and SC32A as were used at the tower both components could be monitored in the lab. Retrieved the tower 10X program from the data logger and reloaded into storage module and data logger. Returned both the original storage module and data logger to the system but again was unable to monitor logger or link with storage module. Removed the SC32A cover and found that the jumper was removed. In the past this has caused intermittent communications problems with different systems. Placed the jumper in the circuit and was then able to successfully monitor the data logger and link with the storage module. System RTS at 1900 GMT. SDS successfully collected a file at 1904 GMT and the raw data was viewed in /data/raw/twr10xC1/in. Data looked good. RTS status O.K.

8/6/98

GRS

Lowered the 60m tower booms on side A at 1920 GMT to read some program values on the 25m IRT. Raised the booms and RTS the instrument at 1939 GMT.

7/16/98

7/17/98

CAM/DC

David Cook on site for 6 month calibrations. Lowered the booms at 1458 GMT on 7/16/98. R/R *Temp/Hum sensor for 6 month calibration. Performed ice bath and sensor checks on new sensor. Left booms connected at ground level overnight. Returned on 7/17/98 and rechecked sensors. Status O.K. Raised booms and RTS at 1505 GMT. *Old - S1720003 *New - S1720002

6/23/98

MW

SDS operations reported a loss of communications with all of the 60M tower instruments beginning at 1004 GMT today. Power cycled the tower COS at 1404 GMT. Communications restored after the power cycle. RTS at 1405 GMT.

5/1/98

CM/MW

The A side 25M and 60M booms were lowered at 1826 GMT for chilled mirror removal. Raised booms and RTS at 1935 GMT.

12/19/97

CAM

Lowered booms from 2100 GMT to 2200 GMT for chilled mirror maintenance.

9/27/97

DKB

IOP reports no files from the 60M T/H. Reset GFCI breaker at 0245 GMT.

9/26/97

CAM

South booms lowered from 2010 to 2047 GMT to R/R 25M T/H.

9/12/97

MW

Jeff Slater reported garbled characters coming back from the instrument. Checked the continuity of all the fibers at 1830 GMT. All fibers OK. Noticed a steady light being emitted from the transmit fiber port of the bit driver located in the instrument box. Disconnected that bit driver from the instrument and replaced its power source with an external power supply to isolate the bit driver. The bit driver continued to emit a beam out of its transmit port. Replaced both bit drivers in the communications loop with model 2505 S bit drivers at 1900 GMT. The receive led went out on both bit drivers and communications with the instrument through the switch was verified by receiving a star back from the CR10X. Also, was able to monitor the instrument for the first time through the fibers from the aerosol trailer. RTS at 1930 GMT.

9/11/97

CM

Mentor requested that the "A" side 60M T/H aspirator shield be replaced with an extender tube and shield. A Vaisala probe adapter also needed to be installed. The 60M "A" boom was already lowered for IOP instrument installation. Installed the adapter and extender tube and located the probe sensor end 1.75 inches up inside the tube. RTS at 2130 GMT. Note that the boom remained at ground level for completion of IOP instrument installation.

9/11/97

MW/CM

SDS ops reported no communications with the T/H10X out at the 60 M tower. Previous T/S efforts indicated that a fiber loop back was successful from the aerosol trailer to the instrument. It was possible to monitor the instrument by hooking up directly to the data logger but communications through the fiber from the aerosol trailer could not be established. Both bit drivers in the communications loop had a solidly lit receive led. Noticed the bit driver in the instrument box was a 2505 S model and the bit driver which was connected to the COS was a 2505 model. Removed the 2505 S model from the instrument box and replaced it with a 2505 model. Was able to receive a star back from the instrument logger through the COS after the bit driver change. However, the bit driver which was connected to the COS still had a solidly lit receive led. Contacted SDS ops and verified communication to the instrument from the SDS. RTS at 2130 GMT.

9/10/97

DN

Lowered the tower from 1915 to 1950 GMT for IOP instrument installation preliminary.

8/13/97

GS

Calibrated the new T/H sensor. Raised boom "A" East on 8/14/97 at 1549 GMT.

8/12/97

GRS

Lowered the 60 meter boom on the "A" side at 1625 GMT to install a new aspirated temp/hum sensor. Installed an aspirator and a *Vaisala temp/hum sensor, model HMP35D, at the 60 meter height. Also installed **CR10X data logger, ***SM716 storage module, and ****PS12 power supply. The sensor will remain down at the ground level until calibrations are complete. Note the 25M level shares the data logger, storage module and power supply. *New - S1720003 **New - 08317 ***New - 5824 ****New - 13799